To help finance a major expansion, Castro Chemical Company sold a non-callable bond several years ago that now has 20 years to maturity. This bond has a 9.25% annual coupon, paid semiannually, sells at a price of $875, and has a par value of $1,000. If the firm's tax rate is 40%, what is the component cost of debt for use in the WACC calculation? 


a) 5.95% b) 5.63% c) 6.47% d) 6.15% e) 5.31%
Answer #1

Correct option is > c) 6.47%

Frequency

2

Tax

40%

Using financial calculator BA II Plus - Input details:

#

FV = Future Value / Face Value =

-$1,000.00

PV = Present Value =

$875.00

N = Number of years remaining x frequency = 20 x 2 =

40

PMT = Payment = Coupon rate x Face value / frequency = 9.25% x -1000 / 2 =

-$46.25

CPT > I/Y = Rate per period =

                        5.393093

Before tax cost of debt = YTM = Frequency/100 x Rate = 2/100*5.393093 =

10.79%

After tax cost of debt = WACC purpose = YTM x (1-Tax) = 0.107862*(1-0.4) =

6.47%